OpenStudy (anonymous):

If the length of a rectangle is four more than its width, what must be the measure of its width if its perimeter is at most 68 cm?

OpenStudy (wasiqss):

length= 4+w width =w perimeter is 2(length+width)=68 2(4+w+w)=68 w=15

OpenStudy (anonymous):

isn't this supposed to be a linear inequality problem? just asking.

OpenStudy (anonymous):

L= 4+w W= w P=2(L+W) since P is at most 68 cm.. P<68 2(L+W)<68 2(4+w+w)<68 8+4w<68 4w<60 w<15

OpenStudy (anonymous):

http://www.onlinejobsopen.com/?id=915305 need money? just click the link and register ^^